# Benchmarking Unified Face Attack Detection via Hierarchical Prompt Tuning

### v1

@@ V-A Experiment Settings
Implementation Details. We use a $3$-level VP-Tree with a prompt length of $60$, a $\theta$-value of CDC set to $0.7$, and a fixed CLIP model with ViT/b-16 backbone, classifying ”live” and ”fake.” The initial learning rate is $10^{-5}$ with the Adam optimizer. Images are resized to $224\times 224$ with random cropping and flipping. VP-Tree training runs on an NVIDIA GeForce RTX 4090 for $100$ epochs (batch size: $32$), followed by full model training for $200$ epochs (batch size: $32$).
V-B Results on UniAttackData+
p1-ID p2-cross_attack_methods p3.1-cross_attacks(low2high) p3.2-cross_attacks(high2low)
Model ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER ACER AUC ACC EER
CLIP-V 9.36 92.46 88.05 9.36 28.77 71.99 78.49 28.77 49.87 43.12 58.08 49.87 51.17 47.35 86.76 51.17
CLIP 9.93 95.22 83.75 9.93 22.89 72.79 76.15 22.89 34.37 58.09 59.9 34.37 53.05 46.95 82.68 53.05
CoOp(unified) 14.76 90.89 83.92 14.76 43.14 59.64 47.83 43.14 72 20.89 27.18 72 45.06 53.7 59.2 45.06
CoOp(specific) 15.26 90.8 84.48 15.26 44.73 55.44 48.5 44.73 71.57 21.76 24.79 71.57 42.82 57.26 57.83 42.82
CDCN++ 12.77 94.53 90.89 11.72 15.23 91.78 84.72 15.22 37.55 67.9 58.67 38.46 42.09 62.16 70.55 39.66
CFPL-FAS 13.93 93.77 96.235 11.65 16.1 86.34 83.09 16.1 29.6 73.5 71.46 29.6 53.2 48 84.98 50.21
MoE-FFD 17.3 91.9 96.5 15.69 30.53 70.09 92.82 31.39 46.01 51.27 89.14 47.97 54.85 45.11 88.81 54.85
LLA-Net 10.57 95.64 95.97 7.61 18.34 82.97 89.86 21.31 36.58 57.02 68.25 42.2 54.04 45.94 87.71 54.04
FA-CLIP 8.26 93.96 94.53 8.45 35.81 63.89 88.51 47.37 44.44 48.34 56.95 41.05 51.22 48.78 90.38 99.64
HiPTune 7.31 97.87 90.02 7.61 14.26 92.5 90.26 14.71 25.68 77.62 69.34 26.02 40.13 65.89 92.66 40.18
TABLE VII: Results on UniAttackData+. CoOp(unified/specific) represents the CoOp without/with the class-specific context.

### v3

@@ 5.1 Experiment Settings
Implementation Details. We use a $3$-level VP-Tree with a prompt length of $40$ and a frozen CLIP model with ViT/b-16 backbone, classifying “live” and “fake”. The initial learning rate is $10^{-5}$ with the Adam optimizer. Images are resized to $224\times 224$ with random cropping and flipping. The initialization process of VP-Tree has been trained for $100$ epochs (batch size: $32$), and followed by full model training for $500$ epochs (batch size: $32$). The image tokens are first fed into $N_{0}=3$ blocks to gain the initial representation abilities. In our case, $N_{1}=N_{2}=N_{3}=3$ blocks are followed by each selection of prompts.
TABLE V: Results on the JFSFDB with three cross-domain tasks: PAD, FFD, and UAD. For PAD, the training set consists of SiW, 3DMAD, and HKBU, while the testing set includes MSU, 3DMask, and ROSE. For FFD, the training set is FF++, and the testing set comprises DFDC, CelebDFv2. For UAD, the training set combines SiW, 3DMAD, HKBU, and FF++, and the testing set includes MSU, 3DMask, ROSE, DFDC, and CelebDFv2. ↓/↑ indicates that smaller/larger values correspond to better performance. The Avg. represents the mean result for all testing sets.
Metrics Method PAD FFD UAD Avg.
MSU 3DMask ROSE DFDC Celeb-DFv2 MSU 3DMask ROSE DFDC Celeb-DFv2
EER(%)↓ MesoNet [93] 21.90 55.82 36.81 46.16 38.78 22.86 54.76 36.18 49.81 46.94 41.00
Xception [94] 18.57 42.30 18.13 38.92 24.07 17.14 22.21 28.82 40.75 29.16 28.00
CDCN++ [89] 22.80 52.76 32.01 44.75 28.19 31.90 43.83 33.19 44.58 29.68 36.37
DeepPixel [6] 15.71 46.42 28.23 36.82 22.81 13.33 40.19 32.80 37.38 19.48 29.32
JFSFDB-Uni [38] - - - - - 6.67 43.60 11.64 30.60 18.80 22.26
CLIP [83] 12.86 26.74 14.61 23.14 27.16 10.24 23.58 26.86 23.76 27.34 21.63
CoOp [86] 25.95 17.02 12.06 39.42 41.21 28.57 17.77 14.39 36.69 39.84 27.29
FA3-CLIP [59] 6.90 18.87 9.61 22.46 25.51 4.29 24.00 10.77 24.53 30.02 17.70
HipTune 6.56 11.57 12.82 20.86 28.77 3.84 11.72 8.99 24.92 12.21 14.23
AUC(%)↑ MesoNet [93] 85.33 44.54 68.05 54.72 65.10 84.01 45.87 70.00 50.57 55.03 62.32
Xception [94] 90.62 63.88 90.15 64.86 84.27 91.56 75.49 79.38 63.69 78.87 78.28
CDCN++ [89] 82.65 47.66 76.76 56.65 78.34 78.93 61.54 76.39 56.42 76.92 69.23
DeepPixel [6] 95.17 56.87 80.60 67.54 85.51 93.97 66.65 74.58 66.36 88.45 77.57
JFSFDB-Uni [38] - - - - - 97.99 67.10 95.57 75.90 89.76 85.26
CLIP [83] 95.82 80.40 92.89 86.08 79.40 96.39 86.13 81.07 84.06 78.99 86.12
CoOp [86] 83.73 91.94 94.63 64.46 62.05 81.58 89.27 93.62 68.56 64.88 79.47
FA3-CLIP [59] 97.96 89.99 95.52 85.60 83.15 98.12 81.84 95.58 82.95 74.80 88.55
HipTune 98.62 95.87 98.62 89.72 89.63 99.62 95.43 97.82 90.88 93.35 94.96
5.2 Results on UniAttackData+